# Saleh Ashkboos ETH Zurich

### v1

@@ Abstract
We introduce QuaRot, a new Quantization scheme based on Rotations, which is able to quantize LLMs end-to-end, including all weights, activations, and KV cache in 4 bits. QuaRot rotates LLMs in a way that removes outliers from the hidden state without changing the output, making quantization easier. This computational invariance is applied to the hidden state (residual) of the LLM, as well as to the activations of the feed-forward components, aspects of the attention mechanism and to the KV cache. The result is a quantized model where all matrix multiplications are performed in 4-bits, without any channels identified for retention in higher precision. Our quantized Llama 2-70B model has losses of at most 0.29 WikiText-2 perplexity and retains 99% of the zero-shot performance. Code is available at: https://github.com/spcl/QuaRot.
@@ 1 Introduction
Large language models (LLMs) have become increasingly important due to their countless applications. However, their inference requires a significant amount of computations, memory, and energy. LLM inference has two main phases: during the prefill phase, the prompt is processed and cached in each layer, and then in the decoding phase the output tokens are generated one-by-one. The prefill phase is usually compute-bound while the decoding phase is known to be memory-bound.
Quantization is among the most important techniques to solve both memory and compute issues during LLM inference. Joint quantization aims to reduce the precision of parameters and KV cache (which results in lower memory usage) as well as inputs (known as activations) and compute the forward pass in low precision. However, quantizing the activations is hard as they have large outlier elements (see Figure 1 for an illustrative example) with much larger values, making activation quantization more difficult than weight quantization, especially for the 4-bit case. Previous work relies on characterizing outlier features using a calibration set and keeping them in higher precision during the inference.
@@
  • Using computational invariance from SliceGPT (Ashkboos et al., 2024), we apply randomized Hadamard transformations to the weight matrices without changing the model. This completely eliminates the outlier features and makes the activations easy to quantize without changing the output of the model.
  • We also apply online Hadamard transformations to the attention module to remove outlier features in keys and values, enabling the KV cache to be quantized.
  • Using the above modifications, QuaRot enables 4-bit LLM inference by quantizing all the weights, activations, and KV caches using integer quantization. On Llama 2-7B model, QuaRot achieves up to 2.16x prefill speedups, and 3.39x memory saving during the decoding stage, with at most 0.63 WikiText-2 perplexity loss. QuaRot preserves 99% of the accuracy of zero-shot tasks. We show that our 6 and 8-bit quantization (weights and activations) is lossless.
@@ 2 Related Work
The majority of quantization schemes focus on compressing LLMs by using weight-only quantization, (Frantar et al., 2022; Dettmers et al., 2023; Lin et al., 2023; Egiazarian et al., 2024; Tseng et al., 2024). These methods downcast each weight into a low precision representation and upcast it before the actual computation. The main computation is still performed in high precision. Several works show that, unlike weights, quantizing the activations is hard due to the outlier features (Wei et al., 2022; Dettmers et al., 2022; Xiao et al., 2023). For 8-bit quantization, LLM.int8() (Dettmers et al., 2022) identifies the outlier features during the inference and keeps them in 16-bits which results in poor performance. SmoothQuant (Xiao et al., 2023) normalizes the features using some scaling factors from a calibration set, solving the issue for the 8-bit case. For 4-bit quantization, recent studies identify and keep the outlier features in high precision offline. Atom (Zhao et al., 2023) keeps 128 outliers in 8-bits and quantizes the inputs using group-wise quantization (through a complex kernel). QUIK (Ashkboos et al., 2023) keeps 256 outlier features with per-token quantization but at the cost of keeping the down-projection layer in 8-bits.
SliceGPT (Ashkboos et al., 2024) is a structured pruning method that produces a model with a smaller hidden dimension. The method relies on computational invariance which means that the hidden state of a transformer model can be computed in any transformed coordinate system, by multiplying input (output) weight matrices of each block on the left (right) by an orthogonal matrix. In this work, we apply computational invariance to LLMs to enable more effective quantization: by applying Hadamard orthogonal matrices to each weight matrix, we change the coordinate system of the hidden state, eliminating outliers. We extend this idea further than SliceGPT, applying coordinate transforms to activations within blocks, and to the KV cache.
@@
Two weight-only quantization methods, QuIP (Chee et al., 2024) and QuIP# (Tseng et al., 2024) have previously considered improving quantization by applying rotations. (Chee et al., 2024) introduced the idea of incoherence processing which applies rotation matrices to the left and right of each weight matrix, as well as the Hessian, which is used in minimizing the weight-quantization objective.
Incoherence processing replaces the weight matrix ${\mathbf{W}}$ with $\tilde{\mathbf{W}}={\mathbf{U}}{\mathbf{W}}{\mathbf{V}}$, where ${\mathbf{U}}$ and ${\mathbf{V}}$ are orthogonal matrices. This suppresses outliers in the weights, making quantization more effective. In the forward pass of the quantized network, the effects of ${\mathbf{U}}$ and ${\mathbf{V}}$ must be undone. To make this fast, QuIP used Kronecker-structured orthogonal matrices which allows these additional computations to be computed quickly. Building on this, QuIP# introduced the use of Hadamard matrices instead of Kronecker-structured matrices. This enables both improved quantization (through better incoherence) and a faster forward pass since the Hadamard transform can be computed in $\mathcal{O}(n\log n)$ addition operations.
@@
Finally, KV cache quantization is another line of research where the goal is to compress the cached keys and values during the generation task. This is crucial for large batch size and long-context length generation as the KV cache will be the main memory bottleneck in such problems. (Sheng et al., 2023) quantizes the KV cache using 4-bit group-wise quantization. KVQuant (Hooper et al., 2024) pushes this limit to 3-bit quantization and KIVI (Liu et al., 2024) shows promising results on 2-bit KV cache quantization. Such methods show that outliers also exist in the keys, and apply a set of complex ideas (like feature-wise quantization, non-uniform representation, keeping high precision outliers) to recover the accuracy of a quantized KV cache.
@@ 3 Background
Here we introduce some background mathematical concepts and notation that are necessary for QuaRot.
@@ 3.1 Orthogonal, Rotation and Hadamard Matrices
An orthogonal matrix ${\mathbf{Q}}$ is a square matrix such that ${\mathbf{Q}}{\mathbf{Q}}^{\top}={\mathbf{I}}$. In this work, we consider only real orthogonal matrices. A rotation matrix is an orthogonal matrix with $|{\mathbf{Q}}|=1$.
A Hadamard matrix is an orthogonal matrix with entries proportional to $\{+1,-1\}$. A Walsh-Hadamard matrix is a square matrix of size $2^{n}$, with
@@
For matrix sizes that are not $2^{n}$, the existence of a Hadamard matrix is not guaranteed. A useful list of known Hadamard matrices is available at (Sloane, 2024). Where we require a Hadamard matrix of size $d\neq 2^{n}$, we factorize $d=2^{n}m$, where $m$ is the size of a known Hadamard matrix. Then we use a Kronecker construction ${\mathbf{H}}_{d}={\mathbf{H}}_{2^{n}}\otimes{\mathbf{H}}_{m}$. This allows computation of ${\mathbf{H}}_{d}{\bm{x}}$ in $\mathcal{O}(d(m+n))$ operations.
Following (Tseng et al., 2024) we make use of randomized Hadamard matrices where convenient. Let ${\bm{s}}$ be a vector containing random draws from $\{+1,-1\}$, and $\tilde{\mathbf{H}}={\mathbf{H}}\,\textrm{diag}({\bm{s}})$. It is straightforward to see that $\tilde{\mathbf{H}}$ is also an orthogonal matrix.
@@ 3.2 Incoherence Processing
The idea of incoherence processing was introduced by (Chee et al., 2024) in the context of weight normalization. We define a weight matrix ${\mathbf{W}}$ to be $\mu$-incoherent if
@@
where max is the element-wise max of the matrix, and $mn$ is the number of elements. A weight matrix that has high incoherence is hard to quantize: the largest element is an outlier relative to the magnitude of the average element. (Chee et al., 2024) showed that multiplying a weight matrix on the left and right by an orthogonal matrix can reduce the incoherence, making matrices easier to quantize. In this work we adopt a similar technique, multiplying weight matrices by orthogonal matrices to improve incoherence, though we add fewer operations to the forward pass, and additionally apply incoherence processing to the activations, enabling improved weight and activation quantization. Figure 1 shows the effect of applying incoherence processing to the activations of Llama- 2 .
@@ 3.3 Transformer structures
Large Language Models are neural networks with repeating attention and feed-forward layers. We introduce our notation through Figures 2 and 3, which show the construction of these blocks. We assume that the construction of the network is “pre-norm”, in that each block is preceded by a LayerNorm or RMSNorm operation. We also assume that the feed-forward network uses a gated architecture, as in Llama- 2 , though our methodology is straightforwardly applied to MLP architectures also.
@@
Figure 2: The gated feed-forward network used in most LMs, including the pre-positioned RMSNorm. The input signal is divided by its norm, and re-scaled by parameters $\alpha$. Two linear blocks, $\textbf{W}_{\textrm{up}}$ and $\textbf{W}_{\textrm{gate}}$ are applied. The activation function $\sigma$ is applied to the gated signal, and the two signals are element-wise multiplied together. The final linear block $\mathbf{W}_{\textrm{down}}$ produces the output signal $\mathbf{Y}$. Before quantization, the operations are performed in single (32 bit) or half (16 bit) precision.
[…]
@@ 3.4 Computational Invariance
The computational invariance theorem [Theorem 1, (Ashkboos et al., 2024)] states that the weights and between-block activations in a transformer can be transformed using an orthogonal matrix with no change to the model output. Here we sketch the main idea. If ${\mathbf{W}}_{\textrm{in}}$ is a weight matrix that appears on the left of a transformer block (i.e. ${\mathbf{W}}_{\textrm{gate}},{\mathbf{W}}_{\textrm{up}}$ in Figure 2, or ${\mathbf{W}}_{k},{\mathbf{W}}_{q},{\mathbf{W}}_{v}$ in Figure 3) then we can multiply on the left by an orthogonal matrix ${\mathbf{Q}}$, and cancel out this effect by multiplying the output matrix (${\mathbf{W}}_{\textrm{down}},{\mathbf{W}}_{\textrm{out}}$) by ${\mathbf{Q}}^{\top}$. This applies despite the fact that RMSNorm is applied between the two blocks, so long as no re-scaling happens in the RMSNorm block (and in practise, we absorb any re-scaling into adjacent weight matrices first). Conceptually, this is because RMSNorm divides the activations by their norm, and applying a rotation ${\mathbf{Q}}$ to the activations does not affect the norm. We have the commutation property
@@ 4 Method
QuaRot consists of two stages. In the first stage, the model weights are manipulated (in full precision), and two additional Hadamard operations are inserted into the model’s forward pass. In the second stage, the weights are quantized using some existing method, and quantization operations are added to the forward pass to enable on-line quantization of the activations. By default, we use GPTQ for quantizing weights, whilst activations are quantized on-the-fly using a simple round-to-nearest scheme. Figures 4 and 5 show updated block diagrams for the forward pass with QuaRot modifications, including updated weight matrices, inserted blocks and the bit-width of weights and activations.
@@
Figure 4: QuaRot applied to a FFN component. The RMSNorm scaling ($\bm{\alpha}$) has been absorbed into the weight matrices. The hidden state $\mathbf{X}$ has been rotated by $\mathbf{Q}$, which is cancelled out by the absorption of $\mathbf{Q}^{\top}$ into the first two weight matrices. All weights are stored in INT4, and all activations immediately before the weights are also quantized to INT4. The result of the matmul between the INT4 weights and INT4 activations on a TensorCore is INT32, which we immediately cast (and scale) to FP16. Whilst the signal is still in FP16, we perform a single on-the-fly Hadamard transform before quantizing and computing a (modified) down-proj, which results in a rotated output $\mathbf{YQ}$.
[…]
@@
We first make use of computational invariance to multiply each weight matrix by an orthogonal matrix. To enable this, the linear parts of LayerNorm or RMSNorm are fused into adjacent weight matrices. Figure 4 shows how the feed-forward block of a transformer is modified by removing the scaling operation from RMSNorm ($\textrm{diag}(\bm{\alpha})$) and absorbing into the subsequent weight matrices. We select a randomized Hadamard matrix with size that matches the hidden dimension of the model and pre- or post-multiply each weight matrix. In Figures 4 and 5 this matrix is denoted ${\mathbf{Q}}$. For example the key-projection weight matrix ${\mathbf{W}}_{k}$ is modified as
@@
With the above weight-modifications in place, we have multiplied many weight matrices on one side by a Hadamard matrix, and changed the computed activations between blocks of the transformer. It still remains to improve the quantize-ability of the activations within each block, which we achieve by inserting on-line Hadamard operations.
We first insert a Hadamard operation into the feed-forward network, before the down-projection matrix. This operation is performed in full precision, and implemented using a fast kernel following (Tseng et al., 2024). This operation is implicitly reversed by fusing a Hadamard matrix into the down-projection matrix of the network: ${\mathbf{W}}_{\textrm{down}}\leftarrow{\mathbf{H}}{\mathbf{W}}_{\textrm{down}}$. Combined with the global matrix ${\mathbf{Q}}$, this means that the down-projection matrix now becomes ${\mathbf{H}}{\mathbf{W}}_{\textrm{down}}{\mathbf{Q}}$ (see Figure 4).
@@
Substituting these modifications into equation 8, we see that the computed result of attention remains unchanged. Since the weights for each head are concatenated in the weight representation, we can equivalently perform a single Kronecker structured multiplication:
@@
Using the method above, we can successfully quantize the value vectors. However, key vectors in the attention module are also known to suffer from outliers . Similar to above, we can use a Hadamard rotation to alleviate this issue, allowing us to have a fully quantized KV-cache. First note that the attention scores ${\mathbf{P}}_{1},\ldots,{\mathbf{P}}_{h}$ are computed as:
${\mathbf{Q}}$ $:=\operatorname{Pos}({\mathbf{X}}{\mathbf{W}}_{q})=:\textrm{concat}[\operatorname{Pos}({\mathbf{Q}}_{1}),\ldots,\operatorname{Pos}({\mathbf{Q}}_{n_{h}})]$ (12)
${\mathbf{K}}$ $:=\operatorname{Pos}({\mathbf{X}}{\mathbf{W}}_{k})=:\textrm{concat}[\operatorname{Pos}({\mathbf{K}}_{1}),\ldots,\operatorname{Pos}({\mathbf{K}}_{n_{h}})]$ (13)
${\mathbf{P}}_{h}$ $:=\operatorname{Softmax}(\alpha\operatorname{Pos}({\mathbf{Q}}_{h})\operatorname{Pos}({\mathbf{K}}_{h}^{T})\odot{\mathbf{M}})\,,$ (14)
where $\bm{\alpha}$ is the Softmax scale usually set to $\frac{1}{\sqrt{d_{h}}}$, ${\mathbf{M}}$ is the attention mask (e.g. causal), and $\operatorname{Pos}$ denotes the positional embedding. Previously, positional embedding was only added before the first layer to the input, in which case $\operatorname{Pos}$ is an identity function. However, recent methods such as RoPE (Su et al., 2021) add position information directly to the key and query vectors.
@@
Since both queries and keys are rotated, the final attention scores ${\mathbf{P}}_{1},\ldots,{\mathbf{P}}_{h}$ remains unchanged. We note that an alternative to the above process is caching the keys before applying the positional encoding. This approach (called Pre-RoPE Caching (Hooper et al., 2024)) needs the inverse rotation to be applied online before applying the positional encoding but removes the need to rotate the query vector. It also adds the overhead of rotating the keys and values for every query. Given that at the time of decoding there is a single query vector and many cached key vectors, we decided to use Post-RoPE caching.
@@
With the weights quantized, we are ready to apply operations to the forward pass that quantize the activations. We deliberately leave the computation of RMSNorm (without scaling) in FP32. We quantize the input of the linear layers using symmetric per-token (rows of the input matrix). During symmetric quantization, the row scales are computed by dividing the maximum absolute value of each token by 7. We then divide each row to its corresponding scale and round the result to its nearest integer. The dequantization is also done by casting the INT32 output of GEMM into FP16, multiply the corresponding scale for the row (from input scales) and column (from weight scales).
@@
Attention is significantly memory bound for longer sequences and larger batch sizes. Having rotated both keys and values, we can successfully quantize the cache into low bit-width. This reduces the number of IO operations needed. We keep the queries in FP16 and use online softmax calculation similar to Flash Attention (Dao et al., 2022). After a segment of the key and value vectors are loaded from the memory, we dequantize them and compute the dot product in FP16.

### v2

@@ Abstract
We introduce QuaRot, a new Quantization scheme based on Rotations, which is able to quantize LLMs end-to-end, including all weights, activations, and KV cache in 4 bits. QuaRot rotates LLMs in a way that removes outliers from the hidden state without changing the output, making quantization easier. This computational invariance is applied to the hidden state (residual) of the LLM, as well as to the activations of the feed-forward components, aspects of the attention mechanism, and to the KV cache. The result is a quantized model where all matrix multiplications are performed in 4 bits, without any channels identified for retention in higher precision. Our 4-bit quantized Llama 2-70B model has losses of at most 0.47 WikiText-2 perplexity and retains 99% of the zero-shot performance. We also show that QuaRot can provide lossless 6 and 8 bit Llama- 2 models without any calibration data using round-to-nearest quantization. Code is available at github.com/spcl/QuaRot.
@@ 1 Introduction
Large language models (LLMs) have become increasingly important due to their countless applications. However, using these models in practice, known as inference, requires a significant amount of computation, memory, and energy, specifically during the prefill phase, in which the model is supposed to process large prompts and cache them in each layer. Quantization is among the most important techniques to improve both memory and compute issues by keeping the data types at lower precision during the forward pass.
As the prefill stage is known to be compute-bound (Ashkboos et al., 2023), joint quantization aims to reduce the precision of parameters and KV cache (which results in lower memory usage) as well as inputs (known as activations) and compute the forward pass in low precision. However, quantizing the activations is hard as they have large outlier elements (see Figure 1 for an illustrative example) with much larger values, making activation quantization more difficult than weight quantization, especially for the 4-bit case. Previous work relies on using a calibration set to characterize the outlier features and keeping them in higher precision for inference (Zhao et al., 2023; Ashkboos et al., 2023).
@@
  • We show that randomized Hadamard transformations can be applied to the weight matrices without additional model modifications. In turn, this completely eliminates outlier features and makes the activations easy to quantize, without changing the output of the model. This can be seen as an extension of the computational invariance idea, proposed in SliceGPT (Ashkboos et al., 2024) in the context of structured pruning.
  • We extend this approach to apply online Hadamard transformations to the attention module to remove outlier features in keys and values, enabling the KV cache to be quantized.
  • Using the above modifications, QuaRot enables 4-bit LLM inference by quantizing all weights, activations, and KV caches using integer quantization. We provide efficient kernel support for QuaRot: on a Llama 2-70B model, QuaRot achieves up to 3.33$\times$ prefill speedups (on a batch size 64 with 2048 sequence length), and 3.89$\times$ memory saving during the decoding stage, with at most 0.47 WikiText-2 perplexity loss. QuaRot preserves 99% of the accuracy of zero-shot tasks and we show that our 6 and 8-bit quantization is lossless with simple round-to-nearest quantization.
@@ 2 Related Work
The majority of quantization schemes focus on compressing LLMs by using weight-only quantization, (Frantar et al., 2022; Dettmers et al., 2023; Lin et al., 2023; Egiazarian et al., 2024; Tseng et al., 2024). These methods downcast each weight into a low-precision representation and upcast it before the actual computation. The main computation is still performed in high precision. Several works show that, unlike weights, quantizing the activations is hard due to the outlier features (Wei et al., 2022; Dettmers et al., 2022; Xiao et al., 2023). For 8-bit case, LLM.int8() (Dettmers et al., 2022) identifies the outlier features during inference and keeps them in 16 bits which results in poor performance. SmoothQuant (Xiao et al., 2023) normalizes the features using some scaling factors from a calibration set, solving the issue for the 8-bit case at the cost of introducing extra hyper-parameters. For 4-bit quantization, recent studies identify the outlier features offline and keep them in high precision. Atom (Zhao et al., 2023) developed a complex kernel for mixed-precision MatMul in the presence of outliers while QUIK (Ashkboos et al., 2023) keeps the down-projection layer in 8 bits.
@@
Two weight-only quantization methods, QuIP (Chee et al., 2024) and QuIP# (Tseng et al., 2024) have previously considered improving quantization by applying rotations. Chee et al. (2024) introduced the idea of incoherence processing which applies rotation matrices to the left and right of each weight matrix, as well as the Hessian, which is used in minimizing the weight-quantization objective. Xi et al. (2023) uses a similar idea during training, using exact Hadamard transformations for each linear layer in the forward pass.
@@
Finally, KV cache quantization is another line of research that aims to compress the cached keys and values during the generation phase. This is crucial for large batch size and long-context length generation as the KV cache will be the main memory bottleneck in such problems. Sheng et al. (2023) quantizes the KV cache using 4-bit group-wise quantization. KVQuant (Hooper et al., 2024) pushes this limit to 3-bit quantization and KIVI (Liu et al., 2024) shows promising results on 2-bit KV cache quantization. Such methods show that outliers also exist in the keys, and apply a set of complex ideas (like feature-wise quantization, non-uniform representation, and keeping high precision outliers) to recover the accuracy of a quantized KV cache.
@@ 3 Background
Here we introduce some mathematical concepts and notation that are necessary for QuaRot.
@@ 3.1 Orthogonal, Rotation and Hadamard Matrices
An orthogonal matrix ${\mathbf{Q}}$ is a square matrix such that ${\mathbf{Q}}{\mathbf{Q}}^{\top}={\mathbf{I}}$. In this work, we consider only real orthogonal matrices. A rotation matrix is an orthogonal matrix. A Hadamard matrix is an orthogonal matrix with entries drawing from $\{+1,\!-1\}$. A Walsh-Hadamard matrix is a square matrix of size $d=2^{n}$, with
@@
For matrix sizes that are not $2^{n}$, the existence of a Hadamard matrix is not guaranteed. A useful list of known Hadamard matrices is made available by Sloane (2024). Where we require a Hadamard matrix of size $d\neq 2^{n}$, we factorize $d=2^{n}m$, where $m$ is the size of a known Hadamard matrix. Then we use a Kronecker construction ${\mathbf{H}}_{d}={\mathbf{H}}_{2^{n}}\otimes{\mathbf{H}}_{m}$. This allows computation of ${\mathbf{H}}_{d}{\bm{x}}$ in $\mathcal{O}(d(m+n))$ operations.
Following Tseng et al. (2024) we make use of randomized Hadamard matrices where convenient. Let ${\bm{s}}$ be a vector containing random draws from $\{+1,\!-1\}$, and $\tilde{\mathbf{H}}={\mathbf{H}}\,\textrm{diag}({\bm{s}})$. It is straightforward to see that $\tilde{\mathbf{H}}$ is also an orthogonal matrix.
@@ 3.2 Incoherence Processing
The idea of incoherence processing was introduced by (Chee et al., 2024) in the context of weight normalization for weight-only LLM quantization. We define a weight matrix ${\mathbf{W}}$ to be $\mu$-incoherent if
@@
where max is the element-wise max of the matrix, and $mn$ is the number of elements. A weight matrix that has high incoherence is hard to quantize: the largest element is an outlier relative to the magnitude of the average element. Chee et al. (2024) showed that multiplying a weight matrix on the left and right by an orthogonal matrix can reduce the incoherence, making matrices easier to quantize. In this work we adopt a similar technique, multiplying weight matrices by orthogonal matrices to improve incoherence, though we add fewer operations to the forward pass. Importantly, we additionally apply incoherence processing to the activations, enabling improved weight and activation quantization. Figure 1 shows the effect of applying incoherence processing to the activations of Llama- 2 .
@@ 3.3 Transformer structures
Large Language Models are neural networks with repeating attention and feed-forward layers. We introduce our notation through Figures 2 and 5, which show the construction of these blocks. We assume that the construction of the network is “pre-norm”, in that each block is preceded by a LayerNorm or RMSNorm operation. We also assume that the feed-forward network uses a gated architecture, as in Llama- 2 , though our methodology is straightforwardly applied to MLP architectures also.
@@
Figure 2: The gated feed-forward network used in most LMs, including the pre-positioned RMSNorm. The input signal is divided by its norm, and re-scaled by parameters $\alpha$. Two linear blocks, $\textbf{W}_{\textrm{up}}$ and $\textbf{W}_{\textrm{gate}}$ are applied. The activation function $\sigma$ is applied to the gated signal, and the two signals are element-wise multiplied together. The final linear block $\mathbf{W}_{\textrm{down}}$ produces the output signal $\mathbf{Y}$. Before quantization, different operations are performed either in single (32 bit) or half (16 bit) precision.
@@ 3.4 Computational Invariance
The computational invariance theorem (Ashkboos et al., 2024, Theorem 1) states that the weights and between-block activations in a transformer can be transformed using an orthogonal matrix with no change to the model output. Here we sketch the main idea. If ${\mathbf{W}}_{\textrm{in}}$ is a weight matrix that appears on the left of a transformer block (i.e., ${\mathbf{W}}_{\textrm{gate}},{\mathbf{W}}_{\textrm{up}}$ in Figure 2, or ${\mathbf{W}}_{k},{\mathbf{W}}_{q},{\mathbf{W}}_{v}$ in Figure 5) then we can multiply on the left by an orthogonal matrix ${\mathbf{Q}}$, and cancel out this effect by multiplying the output matrix (${\mathbf{W}}_{\textrm{down}},{\mathbf{W}}_{\textrm{out}}$) by ${\mathbf{Q}}^{\top}$. This applies despite the fact that RMSNorm is applied between the two blocks, so long as no re-scaling happens in the RMSNorm block (and in practice, we absorb any re-scaling into adjacent weight matrices first). Conceptually, this is because RMSNorm divides the activations by their norm, and applying a rotation ${\mathbf{Q}}$ to the activations does not affect the norm. We have the commutation property
@@ 4 Method
QuaRot consists of two stages. In the first stage, the model weights are manipulated (in full precision), and two additional Hadamard operations are inserted into the model’s forward pass. In the second stage, the weights are quantized using some existing method, and quantization operations are added to the forward pass to enable on-line quantization of the activations (and caches). By default, we use GPTQ (Frantar et al., 2022) for quantizing weights, whilst activations are quantized on-the-fly using a simple round-to-nearest scheme. Figures 3 and 6 show updated block diagrams for the forward pass with QuaRot modifications, including updated weight matrices, inserted blocks and the bit-width of weights and activations.
@@
Figure 3: QuaRot applied to a LLaMa-style FFN. The RMSNorm scaling ($\bm{\alpha}$) has been absorbed into the weight matrices ($(\bm{\alpha})$ is a diagonal matrix with RMSNorm parameters). The hidden state $\mathbf{X}$ has been rotated by $\mathbf{Q}$, which is canceled out by the absorption of $\mathbf{Q}^{\top}$ into the first two weight matrices. All weights are stored in INT4, and all activations immediately before the weights are also quantized to INT4. The result of the matmul between the INT4 weights and activations on a TensorCore is INT32, which we immediately cast (and scale) to FP16 which is the default precision of the model. Whilst the signal is still in FP16, we perform a single on-the-fly Hadamard transform before quantizing and computing a (modified) down-proj, which results in a rotated output $\mathbf{YQ}$.
@@
We first make use of computational invariance to multiply each weight matrix by an orthogonal matrix. To enable this, the linear parts of LayerNorm or RMSNorm are fused into adjacent weight matrices. Figure 3 shows how the feed-forward block of a transformer is modified by removing the scaling operation from RMSNorm ($\textrm{diag}(\bm{\alpha})$) and absorbing into the subsequent weight matrices. We select a randomized Hadamard matrix with size that matches the hidden dimension of the model and pre- or post-multiply each weight matrix. In Figures 3 and 6 this matrix is denoted ${\mathbf{Q}}$. For example the key-projection weight matrix ${\mathbf{W}}_{k}$ is modified as
@@
With the above weight-modifications in place, we have multiplied many weight matrices on one side by a Hadamard matrix and the activations have been changed. It remains to improve the quantization of the activations within each block, which we achieve by inserting on-line Hadamard operations.
We first insert a Hadamard operation into the feed-forward network, before the down-projection matrix. This operation is performed in full precision, and implemented using a fast kernel following Tseng et al. (2024). This operation is implicitly reversed by fusing a Hadamard matrix into the down-projection matrix of the network: ${\mathbf{W}}_{\textrm{down}}\leftarrow{\mathbf{H}}{\mathbf{W}}_{\textrm{down}}$. Combined with the global matrix ${\mathbf{Q}}$, this means that the down-projection matrix now becomes ${\mathbf{H}}{\mathbf{W}}_{\textrm{down}}{\mathbf{Q}}$ (see Figure 3).
@@
Substituting these modifications into equation (8), we see that the computed result of attention remains unchanged. Since the weights for each head are concatenated in the weight representation, we can equivalently perform a single Kronecker structured multiplication:
@@
Using the method above, we can successfully quantize the value vectors. However, key vectors in the attention module are also known to suffer from outliers (Hooper et al., 2024; Liu et al., 2024). Similar to above, we can use a Hadamard rotation to alleviate this issue, allowing us to have a fully quantized KV cache. First note that the attention scores ${\mathbf{P}}_{1},\ldots,{\mathbf{P}}_{h}$ are computed as:
${\mathbf{Q}}$ $\,\leftarrow\,\operatorname{Pos}({\mathbf{X}}{\mathbf{W}}_{q})=\textrm{concat}[\operatorname{Pos}({\mathbf{Q}}_{1}),\ldots,\operatorname{Pos}({\mathbf{Q}}_{n_{h}})]$ (12)
${\mathbf{K}}$ $\,\leftarrow\,\operatorname{Pos}({\mathbf{X}}{\mathbf{W}}_{k})=\textrm{concat}[\operatorname{Pos}({\mathbf{K}}_{1}),\ldots,\operatorname{Pos}({\mathbf{K}}_{n_{h}})]$ (13)
${\mathbf{P}}_{h}$ $\,\leftarrow\,\operatorname{Softmax}(\alpha\operatorname{Pos}({\mathbf{Q}}_{h})\operatorname{Pos}({\mathbf{K}}_{h}^{\top})\odot{\mathbf{M}})\,,$ (14)
where $\bm{\alpha}$ is the Softmax scale usually set to $\frac{1}{\sqrt{d_{h}}}$, ${\mathbf{M}}$ is the attention mask (e.g., causal), and $\operatorname{Pos}$ denotes the positional embedding. Previously, positional embedding was only added before the first layer to the input, in which case $\operatorname{Pos}$ is an identity function. However, recent methods such as RoPE (Su et al., 2021) add position information directly to the key and query vectors.
@@
Since both queries and keys are rotated, the final attention scores ${\mathbf{P}}_{1},\ldots,{\mathbf{P}}_{h}$ remain unchanged. We note that an alternative to the above process is caching the keys before applying the positional encoding. This approach (called Pre-RoPE Caching (Hooper et al., 2024)) needs the inverse rotation to be applied online before applying the positional encoding but removes the need to rotate the query vector. It also adds the overhead of rotating the keys and values for every query. Given that at the time of decoding there is a single query vector and many cached key vectors, we use Post-RoPE caching. This helps us to apply a Hadamard transformation on a single token at each decoding step.
@@
With the weights quantized, we are ready to apply operations to the forward pass that quantize the activations. Following PyTorch implementation, we leave the computation of RMSNorm (without scaling) in FP32. We quantize the input of the linear layers using symmetric per-token (rows of the input matrix). During symmetric quantization, the row scales are computed by dividing the maximum absolute value of each token by 7 (largest representable number in INT4). We then divide each row to its corresponding scale and round the result to its nearest integer. The dequantization is also done by casting the INT32 output of GEMM into FP16, multiply the corresponding scale for the row (from input scales) and column (from weight scales).
@@
Attention is significantly memory bound for longer sequences and larger batch sizes. Having rotated both keys and values, we can successfully quantize the cache into low bit-width. This reduces the number of IO operations needed. We keep the queries in FP16 and use online softmax calculation similar to Flash Attention (Dao et al., 2022). After a segment of the KV vectors are loaded from the memory, we dequantize and compute the dot product in FP16.